OpenStudy (alli6106):

There are 8 sections of seats in an auditorium. Each section contains at least 150 seats but not more than 200 seats. Which of the following could be the number of seats in this auditorium? a.800 b.1000 c.1100 d.1300 e.1700

OpenStudy (mindblast3r):

d

OpenStudy (mindblast3r):

8 sections and 150 minimum seats in a section 150 * 8 = 1200 = minimum number of seats

OpenStudy (alli6106):

how?

OpenStudy (mindblast3r):

max seats = 200 * 8 = 1600

OpenStudy (alexandervonhumboldt2):

no direct answers @Mindblast3r

OpenStudy (alexandervonhumboldt2):

no direct answers @Mindblast3r

OpenStudy (nincompoop):

150 < x < 200

OpenStudy (anonymous):

Mind blaster is correct as if, even if they had 200 seats in each section which is the max number of seats you can have you would end up with 8x200 which is 1600 and the only choice closest to your answer would be the choice D.

OpenStudy (nincompoop):

solve it analytically and properly

OpenStudy (nincompoop):

in a section there is at least 150 seats and not more than 500 seats assign the section as \(x\) and the range to which there are number of possible seats, \(150<x<200 \) then try to look for the least number of seats and maximum number of seats based on the range provided \(8 \times (150 < x < 200) \)

OpenStudy (nincompoop):

there are* at least, I am sorry for the improper grammar in my previous post

OpenStudy (nincompoop):

now it is just a matter of multiplication